OpenStudy (anonymous):

What ticket price maximizes the revenue from ticket sales? Using the following equation: −2500x2−12500x+840000

OpenStudy (anonymous):

First Derive: y' = -5000x-12500 Let x= 0 0= -5000x-12500 Solve for x x=12500/5000 =x x= 2.5 2.5 is the ticket price that maximizes the revenue